### v2

@@ 4 Discussion
We have applied a large collection of machine learning classification techniques to manually collected Russet potato trial datasets. The main goal was to investigate whether and how machine learning methods might provide domain experts with faster and more efficient ways to identify potato varieties for advancement to later trials. We discovered that although missing data can pose a challenge, we can use imputation or strategic dropping of features, as well as engineer new features, to obtain significant improvements in the model classification performances. As illustrated in the results, the large number of negative examples in our dataset resulted in strong performance for predicting when to drop varieties from the trials. Although this bias in the predictions can lead to false negatives, it also means that varieties that will potentially fail in future trials are dropped earlier on, making the process more efficient and economical. Further, our results suggest that integrating machine learning techniques with the art of selection can result in increased selection efficiency. As the variety trials are costly and keeping a clone that is not worthy for one additional year can be cost detrimental. Our results indicate, with availability of extensive data, an integrated art of selection along with data-informed confirmation from machine learning can be effective.
@@ Appendix A Preprocessed Dataset Characteristics
Yield Total in hundredweight/acre. Sum of all different sizes, yield 2’s and culls where each has been multiplied by 8.31. Dataset ranges from 0-1510 with mean of 618. There are 47 missing values.
@@
% No. 1’s Yield No 1’s divided by Yield Total multiplied by 100. Ranges from 0-93. There are 51 missing values.
@@
Yield No. 2’s + $>$ 20oz. Yield total minus Yield No. 1’s minus Yield Under 4 oz. minus Yield culls. Ranges from 0-429. There are 51 missing values.

### v3

@@ 4 Discussion
The integration of machine learning algorithms into the potato breeding program has the potential to significantly impact the selection process and overall efficiency. By leveraging the predictive capabilities of these algorithms, breeders can make more informed decisions regarding which clones to advance to future trials, such as the tri-state trials.
In Oregon state-wide trials, the costs associated with producing and distributing seed tubers, field management, and post-harvest grading can exceed \$2,000 per selection. A significant portion of this expense is attributed to grading, as well as data collection and analysis. Data analysis is also time-intensive, as breeders must dedicate considerable time to meetings and discussions aimed at minimizing biases introduced by data from multiple locations. By automating certain aspects of the selection process, machine learning algorithms such as the ones presented here can save time and resources that would otherwise be spent on manual analysis. Currently, the number of selections tested in two-year Oregon state-wide trials is limited to approximately 50 due to constraints in time and resources. On the other hand, our algorithms enable breeders to assess a larger number of selections in the early stages of trials, effectively addressing bottlenecks. The efficiency gain allows breeders to focus their efforts on more critical tasks, such as evaluating the performance of advanced clones in multi-location trials and making final selection decisions.
The proposed machine learning algorithms, which demonstrated high accuracy in identifying clones to be dropped and moderate performance in predicting clones to be retained, can serve as a valuable tool in the selection process. While these algorithms may not replace the expertise and intuition of experienced breeders, they can complement their decision-making by providing data-driven insights.
By automating certain aspects of the selection process, machine learning algorithms such as the ones presented here can save time and resources that would otherwise be spent on manual analysis. This efficiency gain allows breeders to focus their efforts on more critical tasks, such as evaluating the performance of advanced clones in multi-location trials and making final selection decisions.
Machine learning algorithms can help identify patterns and relationships in the trial data that may not be readily apparent through manual analysis. By considering a wide range of variables and their interactions, the algorithms can provide a more comprehensive assessment of clone performance, taking into account factors such as yield, quality traits, and resistance to environmental stresses.
The simulation study provides valuable insights into the robustness and generalizability of the machine learning models used in the potato breeding program. Lower performance metrics for the simulated datasets in comparison with the original one indicate slight overfitting to the training data. The variability in neural network performance across different distributions emphasizes the need for careful consideration of data characteristics when applying machine learning techniques. The study also highlights the importance of testing models across various data distributions to assess their performance and reliability. By evaluating the models’ performance under different data-generating scenarios, we can better understand their strengths, weaknesses, and potential limitations when applied to real-world scenarios.
However, it is essential to recognize the limitations of machine learning algorithms and computational methods, and emphasize the importance of human expertise in the breeding process. While machine learning can provide valuable predictions and insights, it is ultimately the responsibility of the breeder to interpret the results, consider additional factors, and make final decisions based on their knowledge and experience. We also note that to fully realize the potential of machine learning in the potato breeding program, it is important to continuously refine and validate algorithms using new trial data. Regular updates and improvements to these algorithms can enhance their predictive accuracy and ensure their relevance to the evolving needs of the breeding program.
5 Conclusion
We have applied a large collection of machine learning classification techniques to manually collected Russet potato trial datasets. The main goal was to investigate whether and how machine learning methods might provide domain experts with faster and more efficient ways to identify potato varieties for advancement to later trials.
@@ 5 Conclusion
We discovered that although missing data can pose a challenge, we can use imputation or strategic dropping of features, as well as engineer new features, to obtain significant improvements in the model classification performances. As illustrated in the results, the large number of negative examples in our dataset resulted in strong performance for predicting when to drop varieties from the trials. Although this bias in the predictions can lead to false negatives, it also means that varieties that will potentially fail in future trials are dropped earlier on, making the process more efficient and economical.
Further, our results suggest that integrating machine learning techniques with the art of selection can result in increased selection efficiency. As the variety trials are costly and keeping a clone that is not worthy for one additional year can be cost detrimental. Our results indicate, with availability of extensive data, an integrated art of selection along with data-informed confirmation from machine learning can be effective.
@@ Appendix A Preprocessed Dataset Characteristics
Total Yield in hundredweight/acre. Sum of all different sizes, yield 2’s and culls where each has been multiplied by 8.31. Dataset ranges from 0-1510 with mean of 618. There are 47 missing values.
@@
% No. 1’s Yield No 1’s divided by total yield multiplied by 100. Ranges from 0-93. There are 51 missing values.
@@
Yield No. 2’s + $>$ 20oz. Total yield minus Yield No. 1’s minus Yield Under 4 oz. minus Yield culls. Ranges from 0-429. There are 51 missing values.